### v2

# https://neu-vi.github.io/HOI-Diff/

Abstract
We address the problem of generating realistic 3D human-object interactions (HOIs) driven by textual prompts. To this end, we take a modular design and decompose the complex task into simpler sub-tasks. We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioned on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches. We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt. The APDM is independent of the results by the HOI-DM and thus can correct potential errors by the latter. Moreover, it stochastically generates the contacting points to diversify the generated motions. Finally, we incorporate the estimated contacting points into the classifier-guidance to achieve accurate and close contact between humans and objects. To train and evaluate our approach, we annotate BEHAVE dataset with text descriptions. Experimental results on BEHAVE and OMOMO demonstrate that our approach produces realistic HOIs with various interactions and different types of objects.
Figure 1: HOI-Diff generates realistic motions for 3D human-object interactions given a text prompt and object geometry. Please see the sup. mat. for video results. Darker color indicates later frames in the sequence. Best viewed in color.
1 Introduction
Text-driven synthesis of 3D human-object interactions (HOIs) aims to generate motions for both the human and object that form coherent and semantically meaningful interactions. It enables virtual humans to naturally interact with objects, which has a wide range of applications in AR/VR, video games, and filmmaking, etc.
The generation of natural and physically plausible 3D HOIs involves humans interacting with dynamic objects in various ways according to the text prompts, thereby posing several challenges. First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and object to avoid floating objects. Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g., a person carries a chair, sits on a chair, pushes or pulls a chair). Text-driven 3D HOI synthesis with a diverse set of interactions is still under-explored. Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [11, 34] or textual descriptions [3].
On one hand, recent methods [23, 17, 12, 42, 58, 52, 44, 33] can synthesize realistic human motions for HOIs for static objects only. They usually synthesize the motion in the last mile of interaction, i.e. the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them. On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect real-world complexity. For instance, they focus on grasping small objects [10], provide the object motion as conditioning [25], predict deterministic interactions between the human and the same object without the diversity [54, 37], consider only a small set of interactions (e.g. sit/lift [23], sit/lie down [12], sit [17, 58, 33], grasp [52, 44]), or investigate a single type of object (e.g. chair [17, 58]).
In this paper, we go beyond and propose HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Fig. 1. Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [45] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt. To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt. Our APDM does not rely on the results of the HOI-DM and thus can recover from its potential errors. Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects. Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of the three modules.
To our best knowledge, HOI-Diff is one of the first text-driven approach s capable of generating realistic motions for various HOIs and different types of objects during the interactions. While our approach pioneers this capability, we note the existence of concurrent works such as  [9, 24, 49], which also explore similar directions. For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [3] with text descriptions, which mitigates the issue of severe data scarcity for text-driven 3D HOIs generation. In addition, we evaluate our approach on the OMOMO dataset [25], which focuses on the manipulation of two hands. Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
2 Related Work
Human Motion Generation with Diffusion Models. The denoising diffusion models have been widely used 2D image generations [39, 40, 36] and achieved impressive results. Recent work [56, 45, 5, 18, 38, 1, 2, 4, 6, 41, 43, 46, 51, 55, 57, 59, 53] apply the diffusion model in the task of human motion generation. While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with. Our method is primarily focused on motion generation with human-object interactions.
Scene- and Object-Aware Human Motion Generation. Recent works condition motion synthesis on scene geometry [15, 60, 48, 50]. This facilitates the understanding of human-scene interactions. However, the motion fidelity is compromised due to the lack of paired full scene-motion data. Other approaches [23, 17, 12, 42, 58, 33] instead focus on the interactions with the objects and can produce realistic motions. However, they focus on interacting with static objects with limited interactions. OMOMO [25] proposes a conditional diffusion framework that can generate full-body motion from the object motion. The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion. IMoS [10] synthesizes the full-body human along with the 3D object motions from textual inputs. Although they can generate realistic interactions, IMoS only focuses on grasping small objects with hands. InterDiff [54] predicts whole-body interactions with dynamic objects. However, it focuses on the motion prediction task, where the future motion needs to be forecasted based on past motion information and object information. Note that the interaction type is deterministic, given the object and past motion information. Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt. Recently, there has been a surge of interest in the text-driven synthesis of 3D human-object interactions for dynamic objects, resulting in the development of concurrent works [9, 24, 49].
Multi-Human Motion Generation. Recent works [41, 26, 41] incorporate human-to-human interactions into the motion diffusion process. They usually also adopt cross-attention to exchange information from different human motions. Different from them, our goal is to generate coarse HOIs instead of doing so in a single run, which alleviates the burden of the model.
Affordance Estimation. The affordance estimation on 3D point cloud is studied in [30, 7, 22, 16, 29, 20, 21].  [20] formulates this task as the point cloud segmentation task from the extracted geometric features.  [21] voxelizes the point cloud and creates affordance maps with interactive manipulation.  [22, 16] focuses on the object grasping affordances and some specific applications.  [29] learns the affordance for object-object interactions.  [30] addressed the task of open-vocabulary affordance learning. Overall affordance learning is a very challenging task even the shape of point cloud is given. Instead of predicting the point-wise contact labels, in our method, we propose to make some hypotheses to simplify human-object interactions, making it more tractable without significantly compromising accuracy.
3 Method
3.1 Overview
Motion Representations. We denote a 3D HOI sequence as ${\bm{x}}=\{{\bm{x}}^{h},{\bm{x}}^{o}\}$. It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence. For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [11] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels. For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e., $D^{o}=6$. We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
Diffusion Model for 3D HOI Generation. Given a prompt ${\bm{c}}=({\bm{d}},{\bm{p}})$, consisting of a textual description ${\bm{d}}$ and the object instance’s point cloud ${\bm{p}}$, a diffusion model $p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},{\bm{c}})$ [^1] We use superscripts $h$ and $o$ to denote human and object sequence, respectively. Without a superscript, it means the 3D HOI sequence, containing both ${\bm{x}}^{h}$ and ${\bm{x}}^{o}$. Subscript is used for the diffusion denoising step. learns the reverse diffusion process to generate clean data from a Gaussian noise ${{\bm{x}}_{t}}$ with $T$ consecutive denoising steps
$p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},\bm{c}):=\mathcal{N}({{\bm{x}}_{t-1}},\mathbf{\mu}_{\theta}({{\bm{x}}_{t}},t,\bm{c}),(1-\alpha_{t})\mathbf{I}),$ (1)
where $t$ is the denoising step. Following [45], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models. Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (HOI-DM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) afforance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [31]
${\bm{\mu}}_{t}=\frac{\sqrt{\alpha_{t-1}}\beta_{t}}{1-\alpha_{t}}{\bm{x}}_{0}+\frac{\sqrt{1-\beta_{t}}(1-\alpha_{t-1})}{1-\alpha_{t}}{{\bm{x}}_{t}},$ (2)
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule. $\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [14] is a variance scheduler of choice. Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions. The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions. The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
Figure 4: Illustration of APDM architecture for affordance estimation. Affordance information of human contact labels, object contact positions, and binary object states are represented together as a noise variable, which is fed into the Transformer encoder to generate clean estimation. The object point cloud and textual prompt are taken as conditional input.
Simply adopting the diffusion model described in Eq.(1) would impose a huge burden on the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description. In this paper, we propose HOI-Diff for 3D HOIs generation, disentangling motion generation for humans and objects and estimation of their contacting points. They are later integrated to form coherent and diverse HOIs, which reduces the complexity and burden for each of the three modules, leading to better generation performance as evidenced by our experiments.
Fig. 2 shows the overview of our proposed approach. We introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec. 3.2). To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec. 3.3), whose output will be used as classifier guidance (Sec. 3.4) to correct the interactions at each diffusion step of human/object motion generation.
3.2 Coarse 3D HOIs Generation
First, we introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM) to generate human and object motions that are roughly coherent. As shown in Fig. 4, it consists of two Transformer models [47], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [45]. Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predict clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$. $CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$. It then processes these inputs to generate refined updates based on the cross attention mechanism [47]. The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively. The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [45]. This fine-tuning is critical to ensure the smoothness of the generated human motions. We ablate this design choice in Sec. 4.3. Object MDM is trained from scratch. We modify the input and output linear layers to take in the object motion which has a different dimension from the human motion. More methodology details of HOI-DM can be found in the supplementary material.
3.3 Affordance Estimation
Due to the complexity of the interactions between a human and object, HOI-DM alone usually fails to produce physically plausible results, leading to floating objects or penetrations. To improve the generation of intricate interactions, the problem that needs to be solved is to identify where the contacting areas are between the human and object. InterDiff [54] defines the contacting area based on the distance measurement between the surface of human and object. This approach, however, heavily relies on the quality of the generated human and object motions and cannot recover from errors in the coarse 3D HOI results. In addition, the contact area is diverse even with the same object and interaction type, e.g., “sit” can happen on either side of a table. To this end, we introduce an Affordance Prediction Diffusion Model (APDM) for affordance estimation. As illustrated in Fig. 4, the input includes a text description ${\bm{d}}$ and the object point cloud ${\bm{p}}$. Our APDM doesn’t rely on the results of the HOI-DM and thus can recover from the potential errors in HOI-DM. In addition, it stochastically generates the contacting points to ensure the diversity of the generated motions.
Affordance estimation in 3D point clouds itself is a notably challenging problem [30, 7, 22, 16, 29, 20, 21], especially in the context of 3D HOI generation involving textual prompt. In this paper, we consider eight primary body joints – the pelvis, neck, feet, shoulders, and hands – as the interacting parts in HOI scenarios. It can effectively model common interactions such as grasping an object with both hands, sitting actions involving the pelvis and back, or lifting with a single hand. We use binary contact labels to determine which joints are in contact with the object. Subsequently, we predict eight corresponding contact points on the object surface, identified as the points closest to the selected body joints. Note that the binary contact label estimation for different body joints are independent, allowing us to handle complex HOIs.
Specifically, at each diffusion time step $n$ of APDM [^2] We note that APDM and HOI-DM work independently. We thus use two symbols to denote the different diffusion time steps to avoid confusion., the noisy data consists of human contact labels representing the contact status for the eight primary body joints, denoted as $\bm{y}^{h}_{n}\in\{0,1\}^{8}$, and the eight corresponding contact points on the object surface, denoted as $\bm{y}_{n}^{o}\in\mathbb{R}^{8\times 3}$. The model is designed to predict both contact probabilities and contact positions. Subsequently, dynamic selection of contacting body joints is performed by considering predicted probabilities over a specific threshold $\tau$ (set to be 0.6). The corresponding contact points on the object are then determined based on the selected joints. APDM works similar to the diffusion denoising process described in Eq.(1). Besides, we utilize a large language model (ChatGPT) to determine whether the object state ${\bm{y}}_{0}^{s}\in\{0,1\}$ should be set to static (${\bm{y}}_{0}^{s}=1$) based on the textual description, which can help us better process static objects when synthesizing 3D HOIs, as discussed in the following section. All the clean affordance data is grouped as ${\bm{y}}_{0}=({\bm{y}}^{h}_{0},{\bm{y}}^{o}_{0},{\bm{y}}^{s}_{0})$.
3.4 Affordance-guided Interaction Correction
With the estimated affordance, we can better align human and object motions to form coherent interactions. To this end, we propose to use the classifier guidance [8] to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects.
Algorithm 1 Affordance-guided Interaction Correction 1:Input ${\bm{c}}=({\bm{d}},{\bm{p}})$ consisting of a textual description ${\bm{d}}$ and object point cloud ${\bm{p}}$, HOI-Diff model $\textbf{M}_{\theta}$, objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$, and estimated affordance ${\bm{y}}_{0}=({\bm{y}}_{0}^{h},{\bm{y}}_{0}^{o},{\bm{y}}_{0}^{s})$. 2:${\bm{x}}^{h}_{T},{\bm{x}}^{o}_{T}\leftarrow\text{sample from }\mathcal{N}(\mathbf{0},\mathbf{I})$ 3:$K$ = 1 4:for all $t$ from $T$ to $1$ do 5:     ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}\leftarrow M_{\theta}({\bm{x}}_{t}^{h},{\bm{x}}_{t}^{o},t,{\bm{c}})$ Get ${\bm{\mu}}_{t}^{h},{\bm{\mu}}_{t}^{o}$ according to Eq.(2) with $\Sigma_{t}$ 6:     if $t=1$ then 7:         $K=100$ 8:     end if 9:     for all $k$ from $K$ to $1$ do # Separately perturb 10:         ${\bm{\mu}}^{h}_{t}\leftarrow{\bm{\mu}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}_{t}^{h}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},\bm{y}_{0})$, ${\bm{\mu}}^{o}_{t}\leftarrow{\bm{\mu}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{\mu_{t}^{o}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}_{t}^{o},\bm{y}_{0})$ 11:     end for 12:     $\mathbf{x}_{t-1}^{h}\sim\mathcal{N}({\bm{\mu}}^{h}_{t},\Sigma_{t})$, $\mathbf{x}_{t-1}^{o}\sim\mathcal{N}({\bm{\mu}}^{o}_{t},\Sigma_{t})$ 13:end for 14:return ${\bm{x}}_{0}^{h},{\bm{x}}_{0}^{o}$
Specifically, in a nutshell, we define an analytic function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$ that assesses how closely the generated human joints and object’s 6DoF pose align with a desired objective. In our case, it enforces the contact positions of human and object to be close to each other and their motions are smooth temporally. Based on the gradient of $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$, we can perturb the generated human and object motion at each diffusion step $t$ as in [53, 19, 23],
$\mathbf{{\bm{\mu}}}^{h}_{t}$ $=\mathbf{{\bm{\mu}}}^{h}_{t}-\tau_{1}\Sigma_{t}\nabla_{{\bm{\mu}}^{h}_{t}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0}),$ (3)
$\mathbf{{\bm{\mu}}}^{o}_{t}$ $=\mathbf{{\bm{\mu}}}^{o}_{t}-\tau_{2}\Sigma_{t}\nabla_{{\bm{\mu}}^{o}_{t}}G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0}).$ (4)
Here $\tau_{1}$ and $\tau_{2}$ are different strengths to control the guidance for human and object motion, respectively. Due to the sparseness of object motion features, we assign a larger value to $\tau_{2}$ compared to $\tau_{1}$. This applies greater strength to perturb object motion, facilitating feasible corrections for contacting joints. During the denoising stage, to eliminate diffusion models’ bias that can suppress the guidance signal, we iteratively perturb $K$ times in the last denoising step, as illustrated in Algorithm 1.
How can we define the objective function $G({\bm{\mu}}^{h}_{t},{\bm{\mu}}^{o}_{t},{\bm{y}}_{0})$? We consider three terms here. First, in the generated 3D HOIs, the human and object should be close to each other on the contacting points. We therefore minimize the distance between human contact joints and object contact points
$G_{con}=\sum_{i\in\{1,2,...,8\}}\left\lVert R\big{(}{\bm{\mu}}^{h}_{t}(i)\big{)}-V\big{(}{\bm{y}}^{o}_{t}(i)\big{)}\right\rVert^{2},$ (5)
where ${\bm{\mu}}^{h}_{t}(i)$ and ${\bm{y}}^{o}_{t}(i)$ denote the $i$-th available contacting joint indexed by ${\bm{y}}^{h}_{0}$ and $i$-th object contact point, respectively. $R(\cdot)$ converts the human joint’s local positions to global absolute locations, and $V(\cdot)$ obtains the object’s contact point sequence from the predicted mean of object pose ${\bm{\mu}}^{o}_{t}$.
Second, the generated motion of dynamic objects typically follows human movement. However, we observe that when the human interacts with a static object, such as sitting on a chair, the object appears slightly moved. To address this, we immobilize the object’s movement in the generated samples if the state is static (${\bm{y}}_{0}^{s}=1$), ensuring that proper contact is established between the human and the static object. The objective is defined as
$\small{G_{sta}={\bm{y}}_{0}^{s}\cdot\sum_{l=1}^{L}\left\lVert{\bm{\mu}}_{t}^{o}(l)-\bar{{\bm{\mu}}}_{t}^{o}\right\rVert^{2},}$ (6)
where ${\bm{\mu}}_{t}^{o}(l)$ denotes the object’s 6DoF pose in the $l$-th frame. $\bar{{\bm{\mu}}}_{t}^{o}=\frac{1}{L}\sum_{l}{\bm{\mu}}_{t}^{o}(l)$, which is the average of predicted means of the object’s pose.
Third, we define a smoothness term $G_{smo}(\mu)$ for the object motion to mitigate motion jittering during contact. Due to the space limit, we explain it in the supplementary material.
Finally, we combine all these goal functions to as the final objective
$G=G_{cont}+\alpha G_{sta}+\beta G_{smo},$ (7)
where $\alpha$ and $\beta$ are weights for balance.
4 Experiments
4.1 Setup
Dataset. Since the data designed for studying text-driven 3D HOIs generation is severely scarce, we manually label interaction types, interacting subjects, and contact body parts on top of the BEHAVE dataset  [3]. We then use ChatGPT to assist in generating three text descriptions for each HOI sequence, increasing the diversity of the data. Specifically, BEHAVE encompasses the interactions of 8 subjects with 20 different objects. It provides the human SMPL-H representation [27], the object mesh, as well as its 6DoF pose information in each HOI sequence. To ensure consistency in our approach, we follow the processing method used in HumanML3D [11] to extract representations for 22 body joints. All the models are trained to generate $L=196$ frames in our experiments. In the end, we have 1451 3D HOI sequences along with textual descriptions to train and evaluate our proposed approach. We follow the official train/test split on BEHAVE. We provide more details of the dataset, our annotation process, and annotated textual examples in the supplementary material.
In addition, we evaluate our approach on OMOMO dataset [25]. OMOMO focuses on full-body manipulation with hands. It consists of human-object interaction motion for 15 objects in daily life, with a total duration of approximately 10 hours. It provides text descriptions for each interaction motion. We utilize their object split strategy for both training and evaluation, ensuring the objects between the training and testing sets are different. Additionally, we preprocess human and object motion, similar to our way for the BEHAVE dataset. More details of the OMOMO dataset are in the supplementary material.
Evaluation metrics. We first assess different models for human motion generation using standard metrics as introduced by [11], namely Fréchet Inception Distance (FID), R-Precision, and Diversity. FID quantifies the discrepancy between the distributions of actual and generated motions via a pretrained motion encoder. R-Precision gauges the relevance between generated motions and their corresponding text prompts. Diversity evaluates the range of variation in the generated motions. Additionally, we compute the Foot Skating Ratio to measure the proportion of frames exhibiting foot skid over a threshold (2.5 cm) during ground contact (foot height $<$ 5 cm).
To evaluate the effectiveness of HOIs generation, we report the Contact Distance metric, which quantitatively measures the proximity between the ground-truth human contact joints and the object contact points. Ideally, we should develop similar metrics, e.g., FID, to evaluate the stochastic HOI generation. However, due to the limited data available in BEHAVE [3], training a motion encoder would produce biased evaluation results. To mitigate this issue, we resort to user studies to quantify the effectiveness of different models. Details will be introduced later.
4.2 Comparisons with Existing Methods
[…]
FID $\downarrow$
R-precision $\uparrow$
Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
User Study $\uparrow$
Real 0.096 0.324 7.244 - - -
[…]
MDM* [45] 3.169 0.206 7.606 0.448 0.184 0.87%
PriorMDM* [41] 4.352 0.215 8.186 0.416 0.270 16.53%
InterDiff [54] 7.192 0.170 7.686 0.506 0.218 0.00%
Ours (Full) 1.501 0.235 7.391 0.347 0.182 82.60%
Table 2: Quantitative results on OMOMO dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
FID $\downarrow$
R-precision $\uparrow$
Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
Real 1.539 0.193 4.833 - -
MDM* 9.155 0.137 5.353 0.768 0.191
PriorMDM 11.192 0.128 7.404 0.523 0.344
InterDiff 10.567 0.110 6.493 0.906 0.239
Ours 8.661 0.168 4.780 0.326 0.141
Figure 5: Qualitative comparisons of our approach and baseline methods on BEHAVE dataset. The bottom row, showcasing our method, demonstrates the generation of realistic 3D HOIs with plausible contacts, particularly evident in columns 2 and 4. This contrasts with the baselines, which fail to achieve a similar level of realism and contact plausibility in the interactions. As an additional visual aid, the mesh color gradually darkens over time to represent progression. (Best viewed in color.)
Quantitative Results. Sec. 4.2 reports the quantitative results on BEHAVE dataset [3]. Compared with the baseline methods, our full method achieves the best performance. Specifically, it achieves state-of-the-art results in both FID, R-precision, and Diversity, underscoring its ability to generate high-quality human motions in the context of coherently interacting with objects. The best Contact Distance also suggests that our approach can generate physically plausible HOIs, capturing the intricate interplay interactions between humans and objects.
We also show the results of our User Study in Tab. 4.2 to illustrate user preferences between our method and other baselines. In our user study, we sample 24 sequences from each method and 23 participants are asked to choose their most preferred generation results from these samples. This user study is designed to directly compare user preferences across different methods. The result indicates a strong preference for our method: it is favored over the baselines in a substantial majority of 82.60% of the cases.
Tab. 4.2 presents the quantitative results on the OMOMO dataset. Our method consistently outperforms other baselines by a considerable margin across all metrics. Notably, due to the distinctiveness of objects in the training and testing sets, the results in Tab. 4.2 indicate the effectiveness of our approach in generalizing to unseen objects, proving superior performance compared to other models.
Table 3: Ablation studies of our model’s variants on the BEHAVE dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
FID $\downarrow$
R-precision $\uparrow$
Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
User Study $\uparrow$
w/o Interaction Correction
Ours w/o CM 2.825 0.201 7.850 0.524 0.265 0.00%
Ours w/o pretrain 2.414 0.226 7.915 0.402 0.158 3.90%
[…]
Ours 1.808 0.217 7.718 0.416 0.205 11.69%
w/ Interaction Correction
Ours w/o ${M}^{o}$ CM 3.206 0.205 7.909 0.365 0.310 0.00%
Ours ${}^{Joint}$ 3.412 0.195 7.878 0.421 0.342 -
Ours w/o $G_{con}$ 1.983 0.205 7.826 0.417 0.196 -
Ours w/o $G_{sta}$ 1.636 0.213 7.785 0.367 0.181 -
Ours w/o $G_{smo}$ 1.665 0.205 7.840 0.370 0.182 -
Ours (Full) 1.501 0.235 7.391 0.347 0.182 84.41%
Figure 6: Visual results of different variants of our model in ablation studies on the BEHAVE dataset [3].
Qualitative Results. We showcase qualitative comparisons, rendered with SMPL [27] shapes, between our approach and the baseline methods in Fig. 5. It is observed that the generated HOI motion by other baselines lacks smoothness and realism, where the object may float in the air (e.g., the toolbox in Fig. 5 (b)). Furthermore, these baseline methods struggle to accurately capture the spatial relationships between humans and objects (e.g., the chair in Fig. 5 (e)). In stark contrast, our approach excels in creating visually appealing and realistic HOIs. Notably, it adeptly reflects the intricate details outlined in text descriptions, capturing both the nature of the interactive actions and the specific body parts involved (e.g., raising the trash bin with the right hand in Fig. 5 (a)).